FIND THE EQUATION OF TANGENTS TO CRCLE WITH CENTER AT ORIGIN AND RADIUS 3 WHICH ARE ARE PARALLEL TO THE LINE 3X + 4Y -5=0

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given circle is x2+y2=9; where r=3
slope=m=-3/4
equation of tangent is given by
y=mx\pm r\sqrt{1+m^{2}}
y=-3/4x\pm 3\sqrt{1+-3/4^{2}}
y=(-3x+15)/4
4y=-3x+15
So 3x+4y-15=0 is the equation of tangent
Thanks and REgards
Shaik Aasif
askIITians faculty